Solve a system of four linear equations and four variable:
2x+4y+2z+3w=-7
-3x-2y+5z+4w=11
4x-3y-2z-3w=11
-2x+3y-5z-4w=-16

Answer by Alan3354(69443)

I made an Excel sheet to do these.
I won't learn anything by doing it manually, you might or might not.
Here are the answers so you can check.
w = -531/77
x = 63/77
y = -70/77
z = 604/77